# ST-USleepNet: A Spatial-Temporal Coupling Prominence Network for Multi-Channel Sleep Staging

### v2

@@ 4.1 Datasets
We evaluate ST-USleepNet on three public datasets. ISRUC-S1 [34] contains PSG data from 100 patients (55 male and 45 female), with a total of 87,187 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. ISRUC-S3 [34] contains PSG data from 10 patients (9 male and 1 female), with a total of 8,589 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. MASS-SS3 [35] contains PSG data from 62 healthy people (28 male and 34 female), with a total of 59,304 sleep segments. The PSG data includes 25 channels of EEG, EOG, EMG, and ECG. For each dataset, the PSG data follows the AASM standards [6] and is classified into five categories: Wake, N1, N2, N3, and REM.
@@ 4.2 Experiment Settings
In this study, we use a cross-subject dataset division and perform five-fold cross-validation, ensuring that each data point is used in both the training and testing phases. The evaluation metrics used are accuracy (Acc) and F1 score.
All experiments are performed using NVIDIA A100 Tensor Core GPU. We configure the batch size to 64, utilize the Adam optimizer with a learning rate of 1e-3, and train the model for 80 epochs. Additional information is available in the appendix.
@@ 4.3 Baseline Methods
To comprehensively evaluate the performance of ST-USleepNet, we select 9 models as baselines. These include traditional Random Forest (RF) and Support Vector Machine (SVM) methods [36]; classic CNN and RNN-based methods DeepSleepNet [8] and TinySleepNet [10]; sleep networks based on U-shaped structures U-time [9] and SalientSleepNet [11]; spatial-temporal sleep networks GraphSleepNet [22] and BSTT [33]; and the latest general multivariate time series model FC-STGCN [37].
@@ 4.4 Experiment Results
We compare ST-USleepNet with baseline models in three public datasets. The results are shown in Figure 1, where the best results for each metric are highlighted in bold. Our model achieves state-of-the-art performance across all metrics. This underscores the importance of extracting prominent sleep features and capturing spatial-temporal coupling patterns for sleep staging.
Table 1: Performance comparison with baseline models.
@@
RF 73.81 70.43 73.99 70.20 79.35 77.09
SVM 75.05 72.79 73.56 71.88 81.65 78.23
DeepSleepNet 75.07 69.15 76.70 74.08 80.44 77.02
TinySleepNet 76.92 73.9 76.18 74.20 83.30 77.40
U-time 77.58 76.66 77.70 76.85 80.31 78.36
SalientSleepNet 78.83 78.16 79.34 78.54 85.30 82.35
GraphSleepNet 81.27 79.74 81.82 79.95 88.75 84.42
BSTT 81.96 80.30 82.18 81.54 89.50 85.00
FC-STGNN 77.34 76.77 80.87 78.79 86.23 85.19
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
@@ 4.5 Ablation Study
To verify the effectiveness of the modules designed in ST-USleepNet, we design three variant models by removing the corresponding modules: 1) ST-USleepNet without temporal prominence network (-T). 2) ST-USleepNet without spatial prominence network (-S). 3) ST-USleepNet without spatial-temporal graph construction (-ST). In the ‘-ST’ variant model, we construct a spatial-only graph using the raw EEG signals in the whole segment based on the cosine similarity.
Table 2: Results of ablation study.
@@
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
-T 78.54 77.29 78.63 77.84 87.20 85.77
-4.81% -5.54% -5.19% -5.56% -2.72% -2.52%
-S 78.69 78.04 79.19 77.72 87.94 86.69
-4.63% -4.62% -4.51% -5.70% -1.90% -1.48%
-ST 78.35 77.79 78.52 78.08 87.49 86.07
-5.04% -4.93% -5.32% -5.27% -2.40% -2.18%
@@ 4.6 Hyperparameter Analysis
We explore the impact of various hyperparameters on the model’s performance and identify two with the most significant influence: convolutional kernel/hidden layer arrangement strategies and network depth.
First, ST-USleepNet implements multi-scale strategy in the temporal prominence network by dynamically varying convolutional kernel sizes and in the spatial prominence network by using different hidden layer dimensions, allowing the network to capture features at multiple scales effectively. To validate that the model adopts the optimal strategy, we experimented with the following combinations: 1)increasing then decreasing (our model’s strategy); 2)consistently using small kernels and hidden layers (small); 3)consistently using large kernels (large); 4)decreasing then increasing (reverse). We found that using decreasing convolutional kernels and hidden layer dimensions in the encoder, and increasing them in the decoder, achieves the best performance, as shown in Table 3.
@@
Small 79.76 78.94 79.64 78.29 87.86 85.34
Large 80.16 79.43 80.96 79.46 88.47 86.56
Reverse 78.20 77.65 79.08 77.44 87.29 85.79
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
@@ 5 Conclusion
6 Acknowledgement
This research is supported by A*STAR, CISCO Systems (USA) Pte. Ltd, and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002) and the National University of Singapore President’s Graduate Fellowship.

### v3

@@ 4.1 Datasets
We evaluate ST-USleepNet on three public datasets. ISRUC-S1 Khalighi et al. (2016) contains PSG data from 100 patients (55 male and 45 female), with a total of 87,187 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. ISRUC-S3 Khalighi et al. (2016) contains PSG data from 10 patients (9 male and 1 female), with a total of 8,589 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. MASS-SS3 O’reilly et al. (2014) contains PSG data from 62 healthy people (28 male and 34 female), with a total of 59,304 sleep segments. The PSG data includes 25 channels of EEG, EOG, EMG, and ECG. For each dataset, the PSG data follows the AASM standards Berry et al. (2012) and is classified into five categories: Wake, N1, N2, N3, and REM.
@@ 4.2 Experiment Settings
In this study, we use a cross-subject dataset division and perform 10-fold cross-validation. The evaluation metrics used are accuracy (Acc) and F1-score. All experiments are performed using NVIDIA A100 Tensor Core GPU. We configure the batch size to 64, utilize the Adam optimizer with a learning rate of 1e-3, and train the model for 40 epochs.
@@ 4.3 Baseline Methods
To comprehensively evaluate the performance of ST-USleepNet, we select 7 models as baselines. These include traditional Random Forest (RF) method Dong et al. (2017); classic CNN-based methods DeepSleepNet Supratak et al. (2017) and TinySleepNet Supratak and Guo (2020); sleep networks based on U-shaped structures U-time Perslev et al. (2019) and SalientSleepNet Jia et al. (2021b); spatial-temporal sleep network GraphSleepNet Jia et al. (2020); and the latest general multivariate time series model FC-STGCN Wang et al. (2024b).
@@ 4.4 Experiment Results
The results are shown in Table 1, where the best scores are highlighted in bold. Our model achieves state-of-the-art performance across all metrics on all datasets. The largest performance gains are observed on the ISRUC_S3 dataset. Specifically, ST-USleepNet improves accuracy by +1.14 (1.46%) compared to the strongest baseline Jia et al. (2020), and F1-score by +1.02 (1.31%) compared to the best-performing baseline Jia et al. (2021b). These results underscore the importance of extracting critical features for effective sleep staging.
@@
Random Forest 71.46 70.93 72.19 71.65 80.86 80.54
DeepSleepNet 75.46 75.12 76.83 76.71 82.44 81.56
TinySleepNet 75.87 75.09 77.24 76.93 82.52 82.11
U-time 74.68 74.57 74.36 73.88 82.21 81.83
SalientSleepNet 75.90 75.78 77.10 77.88 84.80 83.73
GraphSleepNet 76.35 76.67 77.97 77.73 84.86 84.27
FC-STGNN 76.29 76.40 76.40 76.85 83.24 82.36
ST-USleepNet 78.28 77.95 79.11 78.90 85.36 84.88
Table 1: Performance Comparison with Baseline Models.
@@ 4.5 Ablation Study
To verify the contribution of the module in ST-USleepNet, we design three variant models by removing the corresponding modules: 1) Without temporal prominence network (-T). 2) Without spatial prominence network (-S). 3) Without spatial-temporal graph construction (-ST). In the ‘-ST’ variant model, we construct a spatial graph using the raw EEG signals in the whole segment based on the cosine similarity.
@@
ST-USleepNet 78.28 77.95 79.11 78.90 85.36 84.88
-T 75.59 74.99 77.00 76.83 83.80 83.19
-3.44% -3.80% -2.67% -2.62% -1.83% -1.99%
-S 75.77 75.41 77.40 77.05 84.07 83.59
-3.21% -3.26% -2.16% -2.34% -1.51% -1.52%
-ST 75.39 75.26 77.38 77.12 83.71 83.23
-3.69% -2.69% -2.19% -2.26% -1.93% -1.94%
Table 2: Results of Ablation Study.
@@ 4.6 Hyperparameter Analysis
We explore the impact of various hyperparameters and identify two with the most significant influence.
First, ST-USleepNet implements a multi-scale strategy in the temporal prominence network by dynamically varying convolutional kernel sizes and in the spatial prominence network by using different hidden layer dimensions, allowing the network to capture features at multiple scales effectively. To validate that the model adopts the optimal strategy, we experimented with the following combinations: 1)increasing then decreasing (our model’s strategy); 2)consistently using small kernels and hidden layers (small); 3)consistently using large kernels (large); 4)decreasing then increasing (reverse). We found that using decreasing convolutional kernels and hidden layer dimensions in the encoder and increasing them in the decoder achieves the best performance, as shown in Table 3.
@@
Small 77.22 77.03 77.93 77.78 84.89 84.72
Large 77.92 77.56 78.64 78.02 84.74 84.09
Reverse 75.87 75.69 76.86 76.53 83.12 82.83
ST-USleepNet 78.28 77.95 79.11 78.90 85.36 84.88
@@ 5 Conclusion
Acknowledgments
This research is supported by A*STAR, CISCO Systems (USA) Pte. Ltd., and the National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002), the National University of Singapore President’s Graduate Fellowship, and the National Natural Science Foundation of China (Grant No. 62306317).